FUTURE VALUE OF AN ANNUITY Find the future values of these ordinary annuities. Compounding occurs once a year. Round your answers to the nearest cent. a. $600 per year for 8 years at 16%. b. $300 per year for 4 years at 8%. c. $600 per year for 12 years at 0%. Rework previous parts assuming that they are annuities due. Round your answers to the nearest cent. d, $600 per year for 8 years at 16%. e. $300 per year for 4 years at 8%. f, $600 per year for 12 years at 0%.
Answer #1

a. PV = 0, N = 8 rate = 16, PMT = 600

use FV function in Excel

future value = 8,544.06

b. PV = 0, N = 4 rate = 8, PMT = 300

future value = 1,351.83

c. PV = 0, N = 12 rate = 0, PMT = 600

future value = 7200

d. future value = 8,544.06*1.16 = 9,911.11

e. FV = 1351.83*1.08 = 1,460.00

f. FV = 7200